Boron is ten years older than his sister. In three years, he will be twice as old as she is now. How old are they now?
Answer by stanbon(75887)

Boron is ten years older than his sister. In three years, he will be twice as old as she is now. How old are they now?
----------------------------
Equations:
B = s + 10
B+3 = 2(s+3)
-----------------------
Substitute for "B" and solve for "s"::
s+10 + 3 = 2s+6
s + 13 = 2s+6
s = 7 (Sister's age)
------
B = s + 10 = 17 (Boron's age now)
-------------------------------------
Cheers,
Stan H.
-------------------